# VLAP: Efficient Video-Language Alignment via Frame Prompting and Distilling for Video Question Answering

### v1

@@ 4.2 Benchmark
VLAP (Ours) 32 72.3 74.9 82.1 75.5 188M -
Table 1: Comparison Results on NExT-QA dataset. Here we measure the accuracy of choosing the right answer. Especially on Temporal and Causal type of questions, our VLAP (using only 4 frames) improves 3.3% and 1.7% respectively, compared with SeViLA. We use bold-face font to indicate the best results and underline on the second best using the same number of frames (brown box for 4 frames and blue box for 8 frames). VLAP using 2-frames only out-performs BLIP-2 using 4-frames by 1.3%. VLAP also achieves upto 3.04$\times$ speedup.
4.3 Baselines
We compare the performance of our VLAP model with several recent prominent models in the field. These include SeViLA [62], BLIP-2 [23], and InternVideo [48], within the context of fine-tuning scenarios. For fair comparison with SeViLA and BLIP-2, we employ the Vision Transformer-Global (ViT-G) as the visual encoder and Flan-T5-XL models as the Large Language Model (LLM). We follow SeViLA to adapt BLIP-2 for video-language training.
5 Results
Here we show both quantitative and qualitative (Sec. 5.2) comparison results between our VLAP and SOTA methods on Video QA and Video Event Prediction task (Sec. 5.1). Together we present an in-depth analysis on the results. Then we conduct ablation study (Sec. 5.3) on our proposed Frame-Prompter and QFormer-Distiller module and the design choices within each module.
5.1 Comparison Results on Video QA and Video Event Prediction
Overall, we demonstrate that the cross-modal key-frame selection matters. Our VLAP model out-performs strong SOTA methods across the 4 Video QA benchmark datasets and 1 Video Event Prediction while keeping the inference latency low. Especially, we highlight that our models’ performance stands out on temporal (including Causal, Interaction, Prediction) type of questions, NExT-QA Temporal (+3.3%, 3.04×3.04$\times$3.04 × speed up), NExT-QA Causal(+1.7%, 1.45×1.45$\times$1.45 × speed up), STAR Interaction (+4.6%, 3.04×3.04$\times$3.04 × speed up), STAR Prediction (+2.8%, 3.04×3.04$\times$3.04 × speed up), How2QA (+0.3%, 3.04×3.04$\times$3.04 × speed up), VLEP (+0.7% with 1.45×1.45$\times$1.45 × speed up, +0.3% with 4.18×4.18$\times$4.18 × speed up) and TVQA (+1.8%, 3.04×3.04$\times$3.04 × speed up).
On the NExT-QA [52] dataset, compared with the SOTA SeViLA [62] on 4-frame and 8-frame setting, our VLAP improves by 1.0% on average accuracy while achieving a 3.04×3.04$\times$3.04 × speed up (see Table 1).
@@ 5.1 Comparison Results on Video QA and Video Event Prediction
Table 2: Comparison Results on STAR Video QA benchmark. For Interaction type of question, our VLAP improved 4.6%. On average, our VLAP out-performs the SOTA method by 2.2% when using 4 frames with 3.04$\times$ speed up. We bold the best results and underline the second best. We’d like to highlight that our VLAP using 2-frames out-performs BLIP-2 using 4-frames.
Method (# Frames) How2QA VLEP TVQA
FrozenBiLM (10)  [55] (NeurIPS2022) 81.5 - 57.5
InternVideo (8)  [48] (Dec 2022) 79.0 63.9 57.2
BLIP-2 (4)  [23] (ICML2023) 82.2 67.0 54.5
SeViLA (4)  [62] (NeurIPS2023) 83.6 68.9 61.6
VLAP (2) (Ours) 82.8 69.2 60.6
VLAP (4) (Ours) 83.9 69.6 63.4
@@
Table 3: Comparison Results on How2QA, VLEP, and TVQA Video QA Benchmarks. VLAP improves performance over SeViLA by 1.8% with 3.04$\times$ speed up on TVQA dataset, 0.7% with 1.45×1.45$\times$1.45 × speed up on VLEP dataset, and 0.3% with 3.04×3.04$\times$3.04 × speed up on How2QA dataset at 4 frames setting. We’d like to highlight that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up.
@@
We also test our VLAP model on the STAR [50] dataset. This dataset is challenging due to its diversified type of questions. As shown in Table 2, our VLAP model outperforms the several strong SOTA models on average by 2.2% with ${\bf 3.04}\times$ speed up. Especially on the STAR Interaction and Prediction type of questions, we outperform SeViLA [62] by 4.6% and 2.8%. This result further demonstrate the importance of key-frame selection. And our model’s advantages in extracting temporal and causal related key-frames.
@@
We further test our model on the longer and larger scale video Video QA benchmark datasets: TVQA [21] and How2QA [27], as shown in Table 3. On the TVQA [21] dataset, our VLAP outperforms the SOTA method by 1.8% at the 4-frame setting. On How2QA, our VLAP improvement is 0.3% with 3.04×3.04$\times$3.04 × speed up compared with SeViLA [62]. This is partially due to the limited 32-frame teacher-QFormer training. On the other hand, compared with Blip-2 [23], VLAP outperforms by 1.7%. This difference again shows the key-frames selected by our VLAP model aligns better with the input question compared with uniform sampling. To test our algorithm’s capabilities, particularly for event prediction, we conduct an additional series of evaluations on VLEP [22] dataset. At the 4-frame setting, VLAP improves over SeViLA [62] by 0.7% with 1.45×1.45$\times$1.45 × speed up. it’s noteworthy that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up (Table 3).
@@
Figure 4: Key-frame Selection Comparison Results (select 4 frames from 32 frames). We compare frames selected by our VLAP compared with that from the SOTA SeViLA [62] method. Across different type of questions, especially the Causal, Temporal type questions, keyframes selected by our network is more relevant and better related to the question.
5.2 Qualitative Results
@@
We qualitatively compare the key-frames selected by our VLAP with the ones from SeViLA [62] on different type of questions. As shown in Figure 4, our network select more question-related key-frames across different question types (including Causal, Temporal and Description or Factual). In the first (col 1 row 1) example in Figure 4, our VLAP locates the frames that visibly show the “car and the dirt”, but the frames selected by SeViLA focuses on the “road”. In the second (col 2 row 1) example, we locate the frames with the “the man not on the bull at the end”, but the frames selected by SeViLA focuses on the “man on the bull at the end”. And in the fourth (col 2 row 2) example, we locate 3 frames with the “four people” according to the question, but non of the frames selected by SeViLA shows “four people”. For Temporal type of questions, our VLAP is also capable of selecting frames that closer to the action specified in the question. In the seventh (col 1 row 4) and eighth (col 2 row 4) example in Figure 4, we locate key-frames around the “the men on stage” and the “black dog” (vs. SeViLA has 2 frames focusing on the “brown dog”).
@@ 5.3 Ablation Study
Here we demonstrate the effectiveness of each proposed modules: the question-aware Frame-Prompter and QFormer-Distiller. We also validate the decoder design choice within the QFormer-Distiller.
@@
5.3.1 Frame-Prompter and QFormer-Distiller Ablation
@@
base+QFD 64.9 68.6 62.2 73.5
base+QFD+FP 66.5 69.6 63.4 74.4
Table 4: Frame-Prompter and QFormer-Distiller Ablation Results. Across all four VideoQA datasets, we observe that both Question-aware Frame-Prompter (FP) and cross-modal QFormer-Distiller (QFD) contribute significantly to our final performance. We highlight that on STAR, adding our QFormer-Distiller improves the accuracy by 2.9%. Our Frame-Prompter (FP) further boost the accuracy by 1.6%.
@@
5.3.2 QFormer-Distiller Decoder Ablation
@@
Frame Prompter T C D Average
@@
FC+LN+FC 69.7 73.1 81.6 74.1
Table 5: QFormer-Distiller Decoder Ablation on NExT-QA. We find that a simple Fully Connected layer (FC) with a Layer Normalization (LN) works best across Temporal (T), Causal (C), Description (D). It is efficient and effective.
@@
6 Conclusion and Future Work
The advances in Large Language Models (LLMs) bring great boost to natural language understandings. New challenge emerges on how to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks. In this work, we present a new VLAP network with a new question-aware Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment. From our experiments, our VLAP out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions. We demonstrate that cross-modal keyframe selection is key to successful video-language alignment task execution. We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.

### v3

@@ 4.1.1 Benchmark:
ViLA+LoRA (Ours) 4 71.4 74.5 80.3 75.1 188M 208
ViLA (Ours) 32 71.8 75.3 82.1 75.6 188M -
Table 1: Comparison Results on NExT-QA dataset. Here we measure the accuracy of choosing the right answer. Especially on Temporal and Causal type of questions, our ViLA (using only 4 frames) improves 3.3% and 1.7% respectively, compared with SeViLA. We use bold-face font to indicate the best results and underline on the second best using the same number of frames (brown box for 4 frames and blue box for 8 frames). ViLA using 2-frames only out-performs BLIP-2 using 4-frames by 1.3%. ViLA also achieves upto 3.04$\times$ speedup. It needs to be noted that our ViLA achieves 75.1% average accuracy with only 4 frames when we finetune LLM with LoRA [20].
4.1.2 Baselines:
We compare the performance of our ViLA model with several recent prominent models in the field. These include SeViLA [65], BLIP-2 [27], and InternVideo [52], within the context of fine-tuning scenarios. For fair comparison with SeViLA and BLIP-2, we employ the Vision Transformer-Global (ViT-G) as the visual encoder and Flan-T5-XL models as the Large Language Model.
4.2 Results
Here we show both quantitative and qualitative (Sec. 4.2.2) comparison results between our ViLA and SOTA methods on Video QA and Video Event Prediction task (Sec. 4.2.1). Together we present an in-depth analysis on the results. Then we conduct ablation study (Sec. 4.3) on our proposed Frame-Prompter and QFormer-Distiller module and the design choices within each module.
4.2.1 Comparison Results on Video QA and Video Event Prediction:
Overall, we demonstrate that the cross-modal key-frame selection matters. Our ViLA model out-performs strong SOTA methods across the 4 Video QA benchmark datasets and 1 Video Event Prediction while keeping the inference latency low. Especially, we highlight that our models’ performance stands out on temporal (including Causal, Interaction, Prediction) type of questions, NExT-QA Temporal (+3.3%, 3.04×3.04$\times$3.04 × speed up), NExT-QA Causal(+1.7%, 1.45×1.45$\times$1.45 × speed up), STAR Interaction (+4.6%, 3.04×3.04$\times$3.04 × speed up), STAR Prediction (+2.8%, 3.04×3.04$\times$3.04 × speed up), How2QA (+0.3%, 3.04×3.04$\times$3.04 × speed up), VLEP (+0.7% with 1.45×1.45$\times$1.45 × speed up, +0.3% with 4.18×4.18$\times$4.18 × speed up) and TVQA (+1.8%, 3.04×3.04$\times$3.04 × speed up).
On the NExT-QA [56] dataset, compared with the SOTA SeViLA [65] on 4-frame and 8-frame setting, our ViLA improves by 1.0% on average accuracy while achieving a ${\bf 3.04}\times$ speed up (see Table 1).
@@ 4.2.1 Comparison Results on Video QA and Video Event Prediction:
Table 2: Comparison Results on STAR Video QA benchmark. For Interaction type of question, our ViLA improved 4.6%. On average, our ViLA out-performs the SOTA method by 2.2% when using 4 frames with 3.04$\times$ speed up. Note that our ViLA using 2-frames out-performs BLIP-2 using 4-frames.
Method Frames Numbers How2QA VLEP TVQA
FrozenBiLM  [59] (NeurIPS2022) 10 81.5 - 57.5
InternVideo  [52] (Dec 2022) 8 79.0 63.9 57.2
BLIP-2  [27] (ICML2023) 4 82.2 67.0 54.5
SeViLA  [65] (NeurIPS2023) 4 83.6 68.9 61.6
ViLA (Ours) 2 82.8 69.2 60.6
ViLA (Ours) 4 83.9 69.6 63.4
@@
Table 3: Comparison Results on How2QA, VLEP, and TVQA Video QA Benchmarks. ViLA improves performance over SeViLA by 1.8% with 3.04$\times$ speed up on TVQA dataset, 0.7% with 1.45×1.45$\times$1.45 × speed up on VLEP dataset, and 0.3% with 3.04×3.04$\times$3.04 × speed up on How2QA dataset at 4 frames setting. Ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up.
@@
We also test our ViLA model on the STAR [54] dataset. This dataset is challenging due to its diversified type of questions. As shown in Table 2, our ViLA model outperforms the several strong SOTA models on average by 2.2% with ${\bf 3.04}\times$ speed up. Especially on the STAR Interaction and Prediction type of questions, we outperform SeViLA [65] by 4.6% and 2.8%. This result further demonstrate the importance of key-frame selection. And our model’s advantages in extracting temporal and causal related key-frames.
@@
We further test our model on the longer and larger-scale Video QA benchmark datasets: TVQA [25] and How2QA [31], as shown in Table 3. On the TVQA [25] dataset, our ViLA outperforms the SOTA method by 1.8% at the 4-frame setting. On How2QA, our ViLA improvement is 0.3% with 3.04×3.04$\times$3.04 × speed up compared with SeViLA [65]. This is partially due to the limited 32-frame teacher-QFormer training. On the other hand, compared with Blip-2 [27], ViLA outperforms by 1.7%. This difference again shows the key-frames selected by our ViLA model aligns better with the input question compared with uniform sampling. To test our algorithm’s capabilities, particularly for event prediction, we conduct an additional series of evaluations on VLEP [26] dataset. At the 4-frame setting, ViLA improves over SeViLA [65] by 0.7% with 1.45×1.45$\times$1.45 × speed up. it’s noteworthy that ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset by 0.3% with ${\bf 4.2}\times$ speed up (Table 3).
@@
Figure 4: Key-frame Selection Comparison Results (select 4 frames from 32 frames). We compare frames selected by our ViLA compared with that from the SOTA SeViLA [65] method. Across different type of questions, especially the Causal, Temporal type questions, keyframes selected by our network is more relevant and better related to the question.
4.2.2 Qualitative Results:
@@
We qualitatively compare the key-frames selected by our ViLA with the ones from SeViLA [65] on different type of questions. As shown in Figure 4, our network select more question-related key-frames across different question types (including Causal, Temporal and Description or Factual). In the first (col 1 row 1) example in Figure 4, our ViLA locates the frames that visibly show the “car and the dirt”, but the frames selected by SeViLA focuses on the “road”. In the second (col 2 row 1) example, we locate the frames with the “the man not on the bull at the end”, but the frames selected by SeViLA focuses on the “man on the bull at the end”. And in the fourth (col 2 row 2) example, we locate 3 frames with the “four people” according to the question, but none of the frames selected by SeViLA shows “four people”. For Temporal type of questions, our ViLA is also capable of selecting frames that are closer to the action specified in the question. In the seventh (col 1 row 4) and eighth (col 2 row 4) example in Figure 4, we locate key-frames around the “the men on stage” and the “black dog” (vs. SeViLA has 2 frames focusing on the “brown dog”).
@@ 4.3 Ablation Study
Here we demonstrate the effectiveness of each proposed module: the text-guided Frame-Prompter and the question-relevant QFormer-Distiller. We also validate the decoder design choice within the QFormer-Distiller.
@@
4.3.1 Frame-Prompter and QFormer-Distiller Ablation:
@@
base+QFormer-Distiller 64.9 68.6 62.2 73.5
base+Frame-Prompter 65.3 68.8 62.7 73.6
base+QFormer-Distiller+Frame-Prompter 66.5 69.6 63.4 74.4
Table 4: Frame-Prompter and QFormer-Distiller Ablation Results. Across all four VideoQA datasets, we observe that both Text-aware Frame-Prompter and cross-modal QFormer-Distiller contribute significantly to our final performance. We highlight that on STAR, adding our QFormer-Distiller improves the accuracy by 2.9%. Our Frame-Prompter further boost the accuracy by 1.6%.
@@
We further explore how QFormer-Distiller (QFD) affect Frame-Prompter (FP). We find our FP+QFD 4-frames(Acc.74.4%) model’s key-frame selection overlaps 92.3% with the key-frame selected by the FP 8-frames(Acc.74.1%) model. Meanwhile, our FP+QFD 4-frames boosts accuracy by 1.8% compared with FP 4-frames model and the key-frame selected overlaps 56.8% with that of the FP 4-frames model. This helps verify our QFD’s ability to both boost performance and enhance the key-frame selection.
4.3.2 QFormer-Distiller Decoder Ablation:
@@
Frame Prompter Decoder Temporal Causal Description Average
@@
FC+LN+GELU+FC 69.7 73.1 81.6 74.1
Table 5: QFormer-Distiller Decoder Ablation on NExT-QA. We find that a simple Fully Connected layer (FC) with a Layer Normalization (LN) works best across Temporal, Causal, Description. It is efficient and effective. GELU is activation function.
@@
4.4 More Discussion
More recent works [24, 65] have shown powerful networks such as LLMs learns most of the data distribution during the pre-training stage. This is one of the major reason why prompt-learning is very effective. Our work is inspired from the prompt-learning. We hope to leverage LLM through proper visual prompting without affecting the generalization ability of the LLM, and this won’t affect LLMs’ original ability on language tasks.
However, for the VQA task itself, optimal performance in this task often necessitates training the Language Model (LLM). Therefore, we conduct an ablation study on NExT-QA that fine-tunes the LLM by using LoRA during the training. Our ViLA achieves 75.1% average accuracy with only 4 frames. This demonstrates that our ViLA has a strong adaptation ability.
Furthermore, to evaluate the generalization of our ViLA, we replace Flan-T5 with llama on NExT-QA, baseline with llama (68.6%) vs. ViLA with llama ( 72.7%), which shows that ViLA can adapt to different LLMs.
5 Conclusion, Limitation and Future Work
How to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks? In this work, we present a new ViLA network with a new text-guided Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment. From our experiments, our ViLA out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions. For the limitation, we don’t have a change to evaluate on LLMs larger than 13b due to the resource constraints. We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.